Staff:

LBTO: Alex (scence),Juan Carlos(AO)

Observer: Shark-vis team(telescope), Shark-nir (remote),INAF remote

Forecast.

forecast0_20231120.png

forecast1-20231120.png

Details.

Power on LBTISX_DX + LuciDX WFS.

Ocam LBTIs OK

Ocam Luci.

two bad quadrants.
  • power cycle bcu2k
  • reset bin1.
Ocam ok

05:40PM

Humidity above limit.

shutter closed.

take darks TV soul-lbtidx and sx to compare.

06:49PM

Still closes humidity above limit

lbtiSX went noo live.
  • change bin from wfs contorl gui
  • OK
08:18PM

Shark-vis called the night.

09:19PM

closed

11:35PM

AO called the night

WFSs power off.

-- JuanCarlosGuerra - 20 Nov 2023
Topic revision: r5 - 21 Nov 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback